The sum of Ann and Fred's age is 28. In seven years time Ann will be twice as old as Fred. How old are they now?
I need to solve this using simultaneous equations.

Found 2 solutions by htmentor, Edwin McCravy:
Answer by htmentor(1343)

A= Ann's age
F= Fred's age
A + F = 28 [sum of their ages]
A + 7 = 2(F + 7) [Ann will be twice Fred's age in 7 years]
So we have two equations and two unknowns
A + F = 28 -> A = 28 - F
substitute this value into the other equation:
28 - F + 7 = 2F + 14
Solve for F:
3F = 21 -> F = 7
Therefore A = 21

Answer by Edwin McCravy(20060)


The other tutor has already done it for you, so all you have to do is
copy it, turn it in, get credit for work that is not yours, and learn
nothing.